10.2 Erosion costs.

Heavenly Cookie Company reports the following annual sales and costs for its current product​ line:

Chocolate

Chip

​Snicker-

doodle

Peanut

Butter

Lemon

Drop

​Cream-

Filled

  Volume

252,000

201,000

141,000

80,000

92,000

  Price

​$0.50

​$0.45

​$0.53

​$0.49

​$0.58

  Cost

​$0.22

​$0.20

​$0.17

​$0.23

​$0.33

Heavenly is thinking of adding Mississippi Mud brownies to the product line. The​ ultra-rich brownies would sell for $0.97 apiece and cost ​$0.75 to produce. The forecasted brownie volume is 221,000 per year. Introduction of​ brownies, however, will reduce cookie sales by 190,500​, with the following drops in sales per​ cookie: 110,000 in chocolate​ chip, 39,000 in​ snickerdoodle, 25,000 in peanut​ butter, 7,000 in lemon​ drop, and 9,500 in​ cream-filled.

What is the erosion cost of introducing the​ brownies?

What is the net change in annual margin if Mississippi Mud brownies are added to the product​ line?
Answer #1

Erosion of chocolate chip = Drop in Sales * ( Price - Cost) = 110,000 * ( 0.5 -0.22) = 30800

Erosion of snickerdoodle= Drop in Sales * ( Price - Cost) = 39,000 * ( 0.45-0.2) = 9750

Erosion of peanut butter= Drop in Sales * ( Price - Cost) = 25,000 * ( 0.53-0.17) = 9000

Erosion of lemon drop= Drop in Sales * ( Price - Cost) = 7000 * ( 0.49-0.23) = 1820

Erosion of cream filled= Drop in Sales * ( Price - Cost) = 9500 * ( 0.58-0.33) = 2375

a) Total erosion costs = 30800+9750+9000+1820+2375 = 53745

b) margin generated through brownie sales = Sales of brownie * ( 0.97-0.75 ) * 221000 = 48620

Net change in annual margin = 53745 - 48620 = 5125

Please give thumbs up. It will help me a lot.